# How Books Tell a History of Statistics in Portugal: Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others

### v1

@@
How Books Tell a History of Statistics in Portugal:
Works of Foreigners, $\boldsymbol{Estrangeirados}$, and Others
Dinis Pestana1,3,4, Rui Santos2,3
1 DEIO, Faculdade de Ciências da Universidade de Lisboa (FCUL)
2 Escola Superior de Tecnologia e Gestão, Instituto Politécnico de Leiria
3 Centro de Estatística e Aplicações (CEA/UL)
4 Instituto de Investigação Científica Bento da Rocha Cabral
@@ 1 Introduction
Sketching the History of Statistics in Portugal naturally requires discussing the influence of foreigners and “$estrangeirados$”, foreign-influenced individuals(1)(1)(1)  In the 17th and 18th centuries, some Portuguese intellectuals, who were exposed to scientific innovations and the Enlightenment philosophy abroad, had a far-reaching influence on the modernization of culture in Portugal. Among these “$Estrangeirados$” (as they are referred to in Portuguese history) are Ribeiro Sanches, Luís António Verney, Francisco Xavier de Oliveira (Cavaleiro d’Oliveira), D. Luís da Cunha (whose political testament recommended Sebastião José de Carvalho e Melo, later Marquês de Pombal, to King D. José), and the Marquês de Pombal himself. The Pombaline reform of the university in 1772 was certainly due to the influential actions of the $estrangeirados$.
@@
The liberal struggles of the 19th century forced many illustrious Portuguese to exile, particularly in the United Kingdom and/or France. Some of these new estrangeirados, namely Solano Constâncio, Alexandre Herculano, and Oliveira Marreca, were important agents for the dissemination of Statistics in Portugal. The politician Passos Manuel, exiled in La Coruña-Plymouth-Belgium-Paris between 1828 and 1832, carried out important reforms during his brief tenure as Minister of the Kingdom in 1836-1837, both in secondary and higher education. One of the consequences of his reform was the creation of a course of Economia Política [Political Economy], whose Professor, Adrião Forjaz de Sampaio, included the teaching of Princípios de Statistica [Statistical Primer] in its program from 1841 onwards. We use the term $estrangeirado$ in a very general sense: estrangeirados will also be much later Zaluar Nunes (first as a student in Paris, then as a political exile in Brazil, due to the diaspora of intellectuals persecuted by Salazarism), Bento Murteira in London, Gustavo de Castro in Paris, Sebastião e Silva in Italy, Dias Agudo, Tiago de Oliveira, and Pedro Braumann in the USA, and later many scholarship holders who pursued doctorates abroad.
, without neglecting the role of “others”. Plagiating Mae West, who when asked what kind of men she preferred, candidly replied: “Personally, I like two types of men — domestic and foreign”, in our opinion many “Others”, who hadn’t the opportunity of extended partnerships abroad, had also a relevant role in the development of Mathematical Statistics in Portugal, and their contributions must be recognised and praised as much as those of the estrangeirados. Notorious examples are António Simões Neto, a man of culture and one of the early enthusiasts of Bayesian ideas in Portugal, who influenced many students at the Faculty of Sciences of Lisbon (namely Dinis Pestana and Maria Antónia Amaral Turkman), Fernando Rosado (to whom we dedicate this sketch of the history of Statistics in Portugal, thanking his contribution to a “Portugaliae Monumenta Statistica” with his fine edition of the Boletim da Sociedade Portuguesa de Estatística [Bulletin of the Portuguese Statistical Society]), Salomé Cabral or Manuela Neves (recently recipient of the Portuguese Statistical Society Life Achievement Award), whose doctoral dissertations were supervised by Tiago de Oliveira, or Rita Vasconcelos, who contributed so much for the progress of the Probability and Statistics group at the University of Madeira.
@@
Concerning the advancements in Mathematical Statistics in the 20th century up to 1974 (Section 3), we highlight how the teaching of Probability, particularly at the Faculdade de Ciências [Faculty of Sciences], produced excellent professors such as Varennes e Mendonça, Zaluar Nunes, or Sebastião e Silva for Instituto Superior de Agronomia (ISA) [Higher Institute of Agronomy], and Leite Pinto in ISEG. These professors, and Bento Murteira, who worked under Maurice Kendall in London, imported the new statistical concepts, particularly those of Pearson (father and son) and Fisher. Fisher’s influence was significant at ISA, but we should not overlook the potential role the Institute of Anthropology in Coimbra could have played due to the hiring of Stevens, a disciple of Fisher, who published a series of remarkable “Methodological Issues” in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of the University of Coimbra]. Finally, we include some informal notes about the last fifty years (Section 4), for which it seems premature to adopt a historical perspective.
@@
The Venetian geographer and statistician Adriano Balbi (1782-1848) was in Portugal in 1820, before settling in Paris from 1821 to 1832, where he changed his name to Adrien Balbi. It was in Paris, in 1822, that he published the two volumes of the Essai statistique sur le royaume de Portugal et d’Algarve [Statistical Essay on the Kingdom of Portugal and Algarve] and also Variétés politico-statistiques de la monarchie portugaise [Various Political-Statistical Observations on the Portuguese Monarchy].
@@
Some statements in Balbi’s books probably irritated the portuguese establishment, since in 1828 Villela da Silva published negative critical observations on Balbi’s works at the Portuguese Royal Press.
On the other hand, Balbi’s work was highly praised by Solano Constâncio (1777-1846), who in 1822 published an analysis of Variétés politico-statistiques sur la monarchie portugaise in the Anais das Ciências, Artes, e Letras [Annals of Sciences, Arts, and Letters], Volume XVII, 84–103.(2)(2)(2)  Pages 212–222 of the 1995 reedition of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], a selection of works by Solano Constâncio [1], published by the Bank of Portugal, edited by José Luís Cardoso, accessible at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf. The pages cited ahead are from this edition. It should be noted, incidentally, that Solano Constâncio, editor of the above mentioned Anais das Ciências, Artes, e Letras, in this critical review refers to the article Considerações sobre a Estatística [Considerations on Statistics] by Cândido Xavier, published in 1820, in Volume X, pp. 134–172.
Solano Constâncio was a remarkable expatriate polygraph, who studied medicine in the United Kingdom between 1791 and 1797, later living in Paris from 1797 to 1799. He returned to Portugal, but in 1807 emigrated again to Paris. From 1822 to 1826 he lived in the United States of America, then returned to Paris, where he died in 1846. Machado de Sousa (1979) [2] and Cardoso (1995) [3] are interesting sources of information about the wide range of interests of this remarkable intellectual.
Regarding Statistics, in the “Preliminary Discourse” of the Anais das Ciências, das Artes e das Letras, Volume I (1818), pp. 1-37, (pp. 81-99 of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy], Solano Constâncio states that
@@
and on p. 176 of the Anais das Ciências, das Artes e das Letras, in the critical review of An inquiry concerning population, etc…, ou investigação acerca da população e da faculdade de multiplicação da espécie humana; obra destinada a refutar a doutrina do ensaio de M. Malthus sobre esse assunto por W. Godwin, Londres, 1820 [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr. Malthus’s essay on this subject by W. Godwin, London, 1820] (Volume XII (1821), 63–104.):
@@
The Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P. Aillaud, Paris, 1838, contains “Considerações sobre a estatística” [“Considerations on Statistics”] (1, 51–64), reproduced on pp. 271–277 of Leituras e Ensaios de Economia Política [Readings and Essays in Political Economy]; we quote from p. 276:
@@
We find it interesting to note that in the 1821 citation the refutation of Malthus’ exponential population growth, concluding that the human species has very little tendency to increase regularly and permanently foreshadows what Verhulst [4, 5, 6] would later formalize, in 1837, with the logistic model; and that, in the 1838 citation, the notable utility of this science, which is now indispensable for […] enlightening the public has a meaning very close to the famous Thomas Carlyle’s [7] statement in Chap. II of Chartism: “A judicious man uses statistics, not to get knowledge, but to save himself from having ignorance foisted upon him.”.
More interesting still is the use in those excerpts of the words “inferences” and “induction”. For Solano Constâncio, Statistics has an inferential component that goes beyond its merely descriptive component. It is perhaps for this reason that he distinguishes “Estatística”  [Statistics] from “Estadística” [Stadistics], etymogically derived from “Estado” (government), the science of governing or dealing with political affairs, more connected to what is termed official statistics.
It was Francisco Solano Constâncio who, in 1836, in his Novo Diccionario Critico e Etymologico da Lingua Portugueza [New Critical and Etymological Dictionary of the Portuguese Language], defined, for the first time in Portuguese, Statistics as the science “that deals with the enumeration of everything that forms the strength of a nation”. He added a list of subjects: population, agriculture, industry, education, public revenues and expenses, military strength, and property distribution. The Novo Diccionario Critico e Etymologico da Lingua Portugueza achieved enduring editorial success, with 11 editions until 1877.
@@
One of the consequences of the 1836 reform by Passos Manuel was the creation of a Chair of Political Economy at the Faculty of Law (resulting from merging the faculties of Laws and of Canons) of the University of Coimbra. Adrião Forjaz de Sampaio included in the program of the chair an informal presentation of Descriptive Statistics, from the perspective of preparing official statistics, heavily influenced by the journal of works of the Société Française de Statistique Universelle (Fontes de Sousa, 2005 [8]; Ferrão, 2006 [9]).
@@
  • Thus, the pioneer of Statistics teaching at the Faculty of Law of the University of Coimbra was Adrião Forjaz de Sampaio (1810-1874). He published Primeiros Elementos da Sciencia da Statistica [First Elements of the Science of Statistics], in which the influence of Quételet is noted. This manual presents the elements of Descriptive Statistics, giving some prominence to averages and dispersions, with rudimentary mathematical formalism. It is interesting to note that in the title, in later editions Forjaz de Sampaio uses “Stadistics” instead of “Statistics”, and that in 1857 José Dias Ferreira (1837-1907), then still a student, published at the University Press an Ensaio sobre os Primeiros Elementos da Theoria da Estadística do Exmo. Senhor Adrião Pereira Forjaz de Sampaio [Essay on the First Elements of the Theory of Statistics by the Honorable Adrião Pereira Forjaz de Sampaio].
@@
In 1887 José Frederico Laranjo published a new program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra, “with the approval of Dr. Manuel Nunes Giraldes”. Citing Bastien (1997), [10] “Note that the Chair of Political Economy, after its creation in 1836, and until 1871, was mostly held by Adrião Forjaz de Sampaio and after that date, and until 1899, by Manuel Nunes Giraldes. Despite the strong relationship of Frederico Laranjo with the Chair of Political Economy, particularly in terms of teaching, program development, publication of teaching manuals, and doctoral thesis examinations […] he was never awarded the respective chair.”. The lessons of J. F. Laranjo in Princípios de Economia Política [Principles of Political Economy] (1891), reprinted in [11] do not contain notes on Statistics.
@@
The dichotomy between Estatística/Estadística [Statistics/Stadistics] continued to divide those interested in this new science(s) in the 19th century. On March 26, 1852, Alexandre Herculano(3)(3)(3)  A remarkable historian, author of an excellent four volumes very comprehensive História de Portugal [History of Portugal] and of História da Origem e Estabelecimento da Inquisição em Portugal. [History of the Origin and Establishment of the Inquisition in Portugal]. He compiled and edited Portugaliae Monumenta Historica, and was also a highly praised romantic novelist author namely of historical novels. (1810-1877) — who was exiled in England and then in France between 1831 and 1832 — proposed to the Royal Academy of Sciences that the section of Economic and Administrative Sciences should “draft the necessary instructions and a series of statistical questions, in harmony with the current state of science, which the Academy should bring to the Government’s attention, seeking to obtain from it an order for civil, ecclesiastical, and military officials, as well as judicial magistrates and elected authorities of any order and denomination, to respond within the scope of their respective actions to the aforementioned questions, with these responses being transmitted to the Academy.”
@@
This interest of Herculano in Statistics was not an isolated episode in his vast array of interests and activities. Two excerpts from Emigration, in Opúsculos, Volume IV [12] show that Herculano maintained his interest in Statistics, appreciating its contribution to the objective understanding of social phenomena:
@@
António de Oliveira Marreca (1805-1889) also went into exile in England and France, returning to Portugal in 1833. He was highly influential in the institutionalization of Political Economy in Portugal, for detailed information cf.  Veríssimo Serrão (1978) [13] and Almodôvar (1995) [14].
@@
  • In Parecer e Memoria sobre a proposta, que apresentou o Sr. Alexandre Herculano: para que a secção de sciencias económicas e administrativas redigisse um Projecto de Estadística [A Review and Memoir on Mr. Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Statistics], Typographia da Academia Real das Sciencias, 1854, Oliveira Marreca returns to the use of the terminology $estadistica$.
Oliveira Marreca was one of the administrators of the Imprensa Nacional de Lisboa [National Press of Lisbon] from 1834 to 1838. Arguably Oliveira Marreca influenced the publication at the Imprensa Nacional de Lisboa of the translation Compendio Estadistico ou Resumo dos Elementos de Estadistica [A Treatise on Statistics, or A Resume of the Elements of Statistics] [15] of the 1847 Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès by A. Moreau de Jonnès (then in charge of the bureau of Statistiques Générales of France). It is worth noting that   Jonnès clearly expresses his understanding of the scope and limitations of Statistics:
@@
For detailed information on Jonnès and his influence, see Paixão Santos (2017) [16].
@@
The politician António José de Ávila (1807-1881) — the presumed inspiration of Eça de Queiroz’s (1925) posthumous satyrical novel O Conde de Abranhos(4)(4)(4)Ávila prohibited a series of liberal conferences (Conferências do Casino), and the outstanding novelist Eça de Queiroz used him as model for the ridiculous politician Conde de Abranhos in a satyrical novel, published only 25 years after the death of Eça de Queiroz. As challenging of the political establishment as Yes, Minister and Yes, Prime Minister. — also reinforced Alexandre Herculano’s proposals in 1854 in a report he prepared for the Minister of Public Works Fontes Pereira de Melo on the works of the Statistical Congress held in Brussels in 1853. He emphasized the importance of establishing uniform bases for national statistical systems, highlighting their importance for the establishment of international comparisons. Cardoso (2012) [17] claims that Ávila “supports the need for a general population census every 10 years, with the collection of basic identification and social status information (profession, education, social and health conditions, etc.). Among the subjects that should be the object of statistical standardization, he gives special attention to the following: territory and property registry; emigration and migratory movements; agricultural production; industrial statistics (production, employment, wages, capital, etc.); commercial statistics; budgets of the working classes; census of the indigent (poverty and social assistance); levels of education; criminality and repression.” Ávila was again the portuguese official delegate at the Statistical Congress held in Berlin in 1863.
@@
There was also a “folkloric” aspect to the use of Statistics: Sexual taboos were also statistically addressed (Santos Cruz, 1841; Azevedo, 1864); and perhaps in an attempt to finance their expenses, some travelers published accounts of their peripatetic wanderings with some pretension of statistical objectivity, such as Portugal a Voo de Pássara(5)(5)(5)  The Flight of a Female Bird in Portugal, as the novelist Camilo Castelo Branco insultingly translated the title Le Portugal à Vol d’Oiseau by Marie Rattazzi, A. Degorce-Cadot, 1880.. In the posthumous publication Londres Maravilhosa [Wonderful London], in the Notes – Excerpts from the notebooks of M. Teixeira-Gomes(6)(6)(6)  Aside being a diplomat and President of the Portuguese Republic, Manuel Teixeira-Gomes was a notorious writer, author namely of the novel Maria Adelaide, published in 1938, describing the attraction for “nymphets” earlier than the famous Lolita of Vladimir Nabokov. Fialho de Almeida was also an important 19th century portuguese novelist, obviously without any serious background in Statistics., he reveals, with a mixture of surprise and irony, that Fialho de Almeida “is preparing for his excursion to Galiza; he plans to make a complete book with statistics and agronomy about Galiza!”
@@
The teaching of elementary Probability Calculus, included in the program of Infinitesimal Calculus, was introduced in 1865 at the Escola Politécnica de Lisboa [Lisbon Polytechnic School], formerly Colégio dos Nobres [Nobleman College] and later, in 1911, Faculdade de Ciências [Faculty of Sciences]. Statistics gradually detached itself from its subordination to Political Economy, and in the last quarter of the 19th century, the first research works linked to Mathematics appeared, authored by Daniel Augusto da Silva (1866, 1870) and Sidónio Paes (1898).

### v2

@@
Abstract
Foreigners and “estrangeirados”, an expression meaning “people going to a foreign country [“estrangeiro”] getting there further education”, had a leading role in the development of Mathematical Statistics in Portugal. In what concerns Statistics, “estrangeirados” in the nineteenth century were mainly liberal intellectuals exiled for political reasons. From 1930 onwards, the research funding authority sent university professors abroad, and hired foreign researchers to stay in Portuguese institutions, and some of them were instrumental in the importation of new concepts and methods of inferential statistics. After 1970, there was a huge program of sending young researchers abroad for doctoral studies. At the same time, many new universities and polytechnic institutes have been created in Portugal. After that, aside from foreigners who choose to have a research career in those institutions and the “estrangeirados” who had returned and created programs of doctoral studies, others, who hadn’t the opportunity of studying abroad, began to play a decisive role in the development of Statistics in Portugal. The publication of handbooks on Probability and Statistics, thesis and core papers in Portuguese scientific journals, and also of works for the layman, reveals how Statistics progressed from descriptive to a mathematical discipline used for inference in all fields of knowledge, from natural sciences to methodology of scientific research.
keywords:
History of Statistics, Portugal, 19th Century, 20th Century.
[…]
@@ 1 Introduction
Sketching the History of Statistics in Portugal naturally requires discussing the influence of foreigners and “$estrangeirados$”, foreign-influenced individuals111  In the 17th and 18th centuries, some Portuguese intellectuals, who were exposed to scientific innovations and the Enlightenment philosophy abroad, had a far-reaching influence on the modernization of culture in Portugal. Among these “$Estrangeirados$” (as they are referred to in Portuguese history) are Ribeiro Sanches, Luís António Verney, Francisco Xavier de Oliveira (Cavaleiro d’Oliveira), D. Luís da Cunha (whose political testament recommended Sebastião José de Carvalho e Melo, later the powerful prime minister Marquês de Pombal, to King D. José), and the Marquês de Pombal himself. The Pombaline reform of the university in 1772 was certainly due to the influential actions of the Estrangeirados. The liberal struggles of the 19th century forced many illustrious Portuguese to exile, particularly in the United Kingdom and/or France. Some of these new estrangeirados, namely Solano Constâncio, Alexandre Herculano, and Oliveira Marreca, were important agents for the dissemination of Statistics in Portugal. The politician Passos Manuel, exiled in La Coruña-Plymouth-Belgium-Paris between 1828 and 1832, carried out important education reforms during his brief tenure as Minister of the Kingdom in 1836-1837, both in secondary and higher education. One of the consequences of his reform was the creation of a course of Economia Política [Political Economy], whose Professor, Adrião Forjaz de Sampaio, included the teaching of Princípios de Statistica [Statistical Primer] [52] in its program from 1841 onwards. We use the term $estrangeirado$ in a very general sense: estrangeirados will also be much later Zaluar Nunes (first as a student in Paris, then as a political exile in Brazil, due to the diaspora of intellectuals persecuted by Salazarism), Bento Murteira in London, Gustavo de Castro in Paris, Sebastião e Silva in Italy, Dias Agudo, Tiago de Oliveira, and Pedro Braumann in the USA, and later many scholarship holders who pursued doctorates abroad., without neglecting the role of “others”. Plagiating Mae West, who when asked what kind of men she preferred, candidly replied: “Personally, I like two types of men — domestic and foreign”, in our opinion many “Others”, who hadn’t the opportunity of extended partnerships abroad, had also a relevant role in the development of Mathematical Statistics in Portugal, and their contributions must be recognised and praised as much as those of the estrangeirados. Notorious examples are António Simões Neto, a man of culture and one of the early enthusiasts of Bayesian ideas in Portugal, who influenced many students at the Faculty of Sciences of Lisbon (namely Maria Antónia Amaral Turkman and Dinis Pestana), Fernando Rosado (to whom we dedicate this sketch of the history of Statistics in Portugal, thanking his contribution to a “Portugaliae Monumenta Statistica” with his fine edition of the Boletim da Sociedade Portuguesa de Estatística [148] [Bulletin of the Portuguese Statistical Society]) and of the Memorial da Sociedade Portuguesa de Estatística [147] [Memorial of the Portuguese Statistical Society], Salomé Cabral or Manuela Neves (recently recipient of the Portuguese Statistical Society Life Achievement Award), whose doctoral dissertations were supervised by Tiago de Oliveira, or Rita Vasconcelos, who contributed so much for the progress of the Probability and Statistics group at the University of Madeira.
@@
Concerning the advancements in Mathematical Statistics in the 20th century up to 1974 (Section 3), we highlight the consequences of starting the teaching of Probability at the Faculdade de Ciências [Faculty of Sciences]. Excellent professors developed the teaching of Statistics at the Instituto Superior de Agronomia (ISA) [Higher Institute of Agronomy] — Varennes e Mendonça, Zaluar Nunes, and Sebastião e Silva — and at the Instituto Superior de Economia e Gestão (ISEG) [Higher Institute of Economy and Management], namely Leite Pinto and Bento Murteira, who worked under Maurice Kendall in London. These professors imported many new statistical concepts, particularly the innovations created by K. Pearson, ‘Student’, Neyman and E. Pearson, Fisher, and their co-workers. Fisher’s influence was significant at ISA, but we should not overlook the potential role the Institute of Anthropology in Coimbra could have played due to the hiring of Stevens, a disciple of Fisher, who published a series of remarkable “Methodological Issues” in the Revista da Faculdade de Ciências da Universidade de Coimbra [Review of the Faculty of Sciences of the University of Coimbra]. Finally, we include some informal notes about the last fifty years (Section 4), for which it seems premature to adopt a historical perspective.
@@
The Venetian geographer and statistician Adriano Balbi (1782-1848) was in Portugal in 1820, before settling in Paris from 1821 to 1832, where he changed his name to Adrien Balbi. It was in Paris, in 1822, that he published the two volumes of the Essai statistique sur le royaume de Portugal et d’Algarve [8] [Statistical Essay on the Kingdom of Portugal and Algarve], and also Variétés politico-statistiques de la monarchie portugaise [9] [Various Political-Statistical Observations on the Portuguese Monarchy].
@@
Some statements in Balbi’s books probably irritated the portuguese establishment, since in 1828 Villela da Silva [204] published at the Portuguese Royal Press negative critical observations on Balbi’s works.
On the other hand, Balbi’s work was highly praised by Solano Constâncio (1777-1846), who in 1822 published an analysis of Variétés politico-statistiques sur la monarchie portugaise [168] in the Anais das Ciências, Artes, e Letras [Annals of Sciences, Arts, and Letters], Volume XVII, 84–103.222  Pages 212–222 of the 1995 reedition of Leituras e Ensaios de Economia Política [171] [Readings and Essays in Political Economy], a selection of works by Solano Constâncio, published by the Bank of Portugal, edited by José Luís Cardoso, accessible at https://www.bportugal.pt/sites/default/files/ocpep-11.pdf. The pages cited ahead are from this edition. It should be noted, incidentally, that Solano Constâncio, editor of the above mentioned Anais das Ciências, Artes, e Letras, in this critical review refers to the article Considerações sobre a Estatística [Considerations on Statistics] by Cândido Xavier [208], published in 1820, in Volume X, pp. 134–172.
Solano Constâncio was a remarkable expatriate polygraph, who studied Medicine in the United Kingdom between 1791 and 1797, later living in Paris from 1797 to 1799. He returned to Portugal, but in 1807 emigrated again to Paris. From 1822 to 1826 he lived in the United States of America, then returned to Paris, where he died in 1846. Machado de Sousa [98] and Cardoso [25] are interesting sources of information about the wide range of interests of this remarkable intellectual.
Regarding Statistics, in the “Preliminary Discourse” [166] of the Anais das Ciências, das Artes e das Letras, Volume I (1818), pp. 1-37, (pp. 81-99 of Leituras e Ensaios de Economia Política [171] [Readings and Essays in Political Economy]), Solano Constâncio states that
@@
and on p. 176 of the Anais das Ciências, das Artes e das Letras, in the critical review of An inquiry concerning population, etc…, ou investigação acerca da população e da faculdade de multiplicação da espécie humana; obra destinada a refutar a doutrina do ensaio de M. Malthus sobre esse assunto por W. Godwin, Londres, 1820 [167] [An inquiry concerning population, etc…, or an investigation into the population and the ability to multiply the human species; a work intended to refute the doctrine of Mr. Malthus’s essay on this subject by W. Godwin, London, 1820] (Volume XII (1821), 63–104.):
@@
The Armazém de Conhecimentos Úteis, nas Artes e Ofícios, ou Colecção de Tratados, Receitas e Invenções de Utilidade Geral [Repository of Useful Knowledge, in Arts and Crafts, or Collection of Treatises, Recipes, and Inventions of General Utility], J.-P. Aillaud, Paris, 1838, contains “Considerações sobre a estatística” [170] [“Considerations on Statistics”] (1, 51–64), reproduced on pp. 271–277 of Leituras e Ensaios de Economia Política[171] [Readings and Essays in Political Economy]; we quote from p. 276:
@@
We find it interesting to note that in the 1821 citation [167] the refutation of Malthus’ exponential population growth, concluding that “the human species has very little tendency to increase regularly and permanently” foreshadows what Verhulst [198, 199, 200] would later formalize, in 1837, with the logistic model; and that, in the 1838 citation [170], “the notable utility of this science, which is now indispensable for […] enlightening the public” has a meaning very close to the famous Thomas Carlyle’s [27] statement in Chap. II of Chartism: “A judicious man uses statistics, not to get knowledge, but to save himself from having ignorance foisted upon him.”.
More interesting still is the use in those excerpts of the words “inferences” and “induction”. For Solano Constâncio, Statistics has an inferential component that goes beyond its merely descriptive component. It is perhaps for this reason that he distinguishes “Estatística”  [Statistics] from “Estadística” [Stadistics], etymologically derived from “Estado” (government), the science of governing or dealing with political affairs, more connected to what is termed official statistics.
It was Francisco Solano Constâncio who, in 1836, in his Novo Diccionario Critico e Etymologico da Lingua Portugueza [169] [New Critical and Etymological Dictionary of the Portuguese Language], defined, for the first time in Portuguese, Statistics as the science “that deals with the enumeration of everything that forms the strength of a nation”. He added a list of subjects: population, agriculture, industry, education, public revenues and expenses, military strength, and property distribution. The Novo Diccionario Critico e Etymologico da Lingua Portugueza achieved enduring editorial success, with 11 editions until 1877.
@@
One of the consequences of the 1836 education reform by Passos Manuel was the creation of a Chair of Political Economy at the Faculty of Law (resulting from merging the faculties of Laws and of Canons) of the University of Coimbra. Adrião Forjaz de Sampaio included in the program of the chair an informal presentation of Descriptive Statistics, from the perspective of preparing official statistics, heavily influenced by the journal of works of the Société Française de Statistique Universelle (Fontes de Sousa, 2005 [51]; Ferrão, 2006 [49]).
@@
  • Thus, the pioneer of Statistics teaching at the Faculty of Law of the University of Coimbra was Adrião Forjaz de Sampaio (1810-1874). He published Primeiros Elementos da Sciencia da Statistica [52] [First Elements of the Science of Statistics], in which the influence of Quételet is noted. This manual presents the elements of Descriptive Statistics, giving some prominence to averages and dispersions, with rudimentary mathematical formalism. It is interesting to note that in the title, in later editions Forjaz de Sampaio uses “Stadistics” instead of “Statistics”, and that in 1857 José Dias Ferreira (1837-1907), then still a student, published at the University Press an Ensaio sobre os Primeiros Elementos da Theoria da Estadística do Exmo. Senhor Adrião Pereira Forjaz de Sampaio [47] [Essay on the First Elements of the Theory of Stadistics by the Honorable Adrião Pereira Forjaz de Sampaio].
@@
In 1887 José Frederico Laranjo [88] published a new program for the Chair of Political Economy and Statistics at the Faculty of Law of the University of Coimbra, “with the approval of Dr. Manuel Nunes Giraldes”. Citing Bastien [15]Note that the Chair of Political Economy, after its creation in 1836, and until 1871, was mostly held by Adrião Forjaz de Sampaio and after that date, and until 1899, by Manuel Nunes Giraldes. Despite the strong relationship of Frederico Laranjo with the Chair of Political Economy, particularly in terms of teaching, program development, publication of teaching manuals, and doctoral thesis examinations […] he was never awarded the respective chair.”. The lessons of J. F. Laranjo in Princípios de Economia Política (1891) [Principles of Political Economy], reprinted in [89], do not contain notes on Statistics.
@@
The dichotomy between Estatística/Estadística [Statistics/Stadistics] continued to divide those interested in this new science(s) in the 19th century. On March 26, 1852, Alexandre Herculano333  A remarkable historian, author of an excellent four volumes very comprehensive História de Portugal [History of Portugal] and of História da Origem e Estabelecimento da Inquisição em Portugal. [History of the Origin and Establishment of the Inquisition in Portugal]. He compiled and edited Portugaliae Monumenta Historica, and was also a highly praised romantic novelist author namely of historical novels. (1810-1877) — who was exiled in England and then in France between 1831 and 1832 — proposed to the Lisbon Royal Academy of Sciences that the section of Economic and Administrative Sciences should “draft the necessary instructions and a series of statistical questions, in harmony with the current state of science, which the Academy should bring to the Government’s attention, seeking to obtain from it an order for civil, ecclesiastical, and military officials, as well as judicial magistrates and elected authorities of any order and denomination, to respond within the scope of their respective actions to the aforementioned questions, with these responses being transmitted to the Academy.
@@
This interest of Herculano in Statistics was not an isolated episode in his vast array of interests and activities. Two excerpts from Emigração [Emigration], in Opúsculos, Volume IV [79] show that Herculano maintained his interest in Statistics, appreciating its contribution to the objective understanding of social phenomena:
@@
António de Oliveira Marreca (1805-1889) also went into exile in England and France, returning to Portugal in 1833. He was highly influential in the institutionalization of Political Economy in Portugal, for detailed information cf.  Veríssimo Serrão [201] and Almodôvar [3].
@@
  • In Parecer e Memoria sobre a proposta, que apresentou o Sr. Alexandre Herculano: para que a secção de sciencias económicas e administrativas redigisse um Projecto de Estadística [103] [A Review and Memoir on Mr. Alexandre Herculano proposal, requiring that the section of economic and administrative sciences establishes a Project on Stadistics], Oliveira Marreca returns to the use of the terminology Estadística.
Oliveira Marreca was one of the administrators of the Imprensa Nacional de Lisboa [National Press of Lisbon] from 1834 to 1838. Arguably Oliveira Marreca influenced the publication at the Imprensa Nacional de Lisboa of the translation Compendio Estadistico ou Resumo dos Elementos de Estadistica [112] [A Treatise on Statistics, or A Resume of the Elements of Statistics] of the 1847 Éléments de Statistique: comprenant les principes généraux de cette science, et un aperçu historique de ses progrès [111] by A. Moreau de Jonnès (then in charge of the bureau of Statistiques Générales of France). It is worth noting that   Jonnès clearly expresses his understanding of the scope and limitations of Statistics:
@@
For detailed information on Jonnès and his influence, see Paixão Santos [130].
@@
The politician António José de Ávila (1807-1881) — the presumed inspiration of Eça de Queiroz’s (1925) posthumous satyrical novel O Conde de Abranhos444Ávila prohibited a series of liberal conferences (Conferências do Casino), and the outstanding novelist Eça de Queiroz used him as model for the ridiculous politician Conde de Abranhos in a satyrical novel, published only 25 years after the death of Eça de Queiroz. As challenging of the political establishment as Yes, Minister and Yes, Prime Minister. — also reinforced Alexandre Herculano’s proposals in 1854 in a report he prepared for the Minister of Public Works Fontes Pereira de Melo on the works of the Statistical Congress held in Brussels in 1853. He emphasized the importance of establishing uniform bases for national statistical systems, highlighting their importance for the establishment of international comparisons. Cardoso [26] claims that Ávila “supports the need for a general population census every 10 years, with the collection of basic identification and social status information (profession, education, social and health conditions, etc.). Among the subjects that should be the object of statistical standardization, he gives special attention to the following: territory and property registry; emigration and migratory movements; agricultural production; industrial statistics (production, employment, wages, capital, etc.); commercial statistics; budgets of the working classes; census of the indigent (poverty and social assistance); levels of education; criminality and repression.” Ávila was again the portuguese official delegate at the Statistical Congress held in Berlin in 1863 [6].
@@
There was also a “folkloric” aspect to the use of Statistics: Sexual taboos were also statistically addressed (Santos Cruz [156]; Azevedo [7]); and perhaps in an attempt to finance their expenses, some travellers published accounts of their peripatetic wanderings with some pretension of statistical objectivity, such as Portugal a Voo de Pássara [142]555  As the novelist Camilo Castelo Branco insultingly translated the title Le Portugal à Vol d’Oiseau [The Flight of a Female Bird in Portugal,] by Marie Rattazzi, A. Degorce-Cadot, 1880. “Pássara” means literally “female bird”, but its slang meaning is “cunning”, “smart-ass”, “shrewed-ass”, or even worse.. In the posthumous publication Londres Maravilhosa [Wonderful London], in the Notes – Excerpts from the notebooks of M. Teixeira-Gomes666  Aside being a diplomat and President of the Portuguese Republic, Manuel Teixeira-Gomes was a notorious writer, author namely of the novel Maria Adelaide, published in 1938, describing the attraction for “nymphets” earlier than the famous Lolita of Vladimir Nabokov. Fialho de Almeida was also an important 19th century portuguese novelist, obviously without any serious background in Statistics. In 1905 he published Cadernos de Viagem. Galiza [Travel Notebooks. Galiza]., he reveals, with a mixture of surprise and irony, that the novelist and journalist Fialho de Almeida “is preparing for his excursion to Galiza; he plans to make a complete book with statistics and agronomy about Galiza!
@@
The teaching of elementary Probability Calculus, included in the program of Infinitesimal Calculus, was introduced in 1865 at the Escola Politécnica de Lisboa [Lisbon Polytechnic School], formerly Colégio dos Nobres [Nobleman College] and later, in 1911, Faculdade de Ciências [Faculty of Sciences]. Statistics gradually detached itself from its subordination to Political Economy, and in the last quarter of the 19th century, the first research works linked to Mathematics were published in 1866 and 1870 (Daniel Augusto da Silva [160, 161]) and later, in 1898, the first doctoral thesis on Mathematical Statistics (Sidónio Paes [129]).